Suppose that a fair die is rolled n times. We say that there is a repeat at the i’th place if the same number occurs on both the i’th and i + 1’st roll. Let X be a random variable representing the number of repeats. Find E[X].
Answer #1

Any query then comment below...i will ecplain you...

Understand carefully...

We take an exple to understand this...

Let we throw 3 times okk..

And repeat at 1st place means we have 1st and 2nd entry same....

So , means on first and secone place we have (1,1) , (2,2) ..... (6,6) ...ok....and 3rd place we have 6 choices...so we have probability is (6*6)/216 = 36/216 = 1/6 ...

Same think happen in n also...  

2 n(nti) 12